OpenStudy (anonymous):

the time spent by customers at a store is: t = 2d^2/3 + 3 where d is the amount of money spent by the customer. to find the mean amount of time spent, do i have to use all different values for money spent, which are provided earlier, or can i use the mean value of the money spent, calculated earlier, to find this out?? pls help :P

OpenStudy (anonymous):

d^2 = (3t-9)/2

OpenStudy (anonymous):

nonononnon, it is "d to the power of 0.66666666666667" or "d to the power of 2/3"

OpenStudy (anonymous):

you cannot use mean of money, as we are not using money, but using money^2/3. for n number of observations, mean time = [2*d1^(2/3) + 3] + [2*d2^(2/3) + 3] + .....+[2*dn^(2/3) + 3] =2*(d1^(2/3)+d2^(2/3)+....+dn^(2/3)) + 3*n

OpenStudy (anonymous):

above is sum of time for mean of time , [2*(d1^(2/3)+d2^(2/3)+....+dn^(2/3)) + 3*n]/n

OpenStudy (anonymous):

tanvi u cud have done that if there was no +3 in r.h.s but as it is there u will get a different answer if u take mean of all the money when compared to taking mean of time and in this case u have to take mean of the time

OpenStudy (anonymous):

so, i WILL have to do the whole procedure, then. thanks :)